# SPICE[Uncaptioned image]: An Automated SWE-Bench Labeling Pipeline for Issue Clarity, Test Coverage, and Effort Estimation

### v1

@@ Abstract
High-quality labeled datasets are crucial for training and evaluating foundation models in software engineering, but creating them is often prohibitively expensive and labor-intensive. We introduce SPICE, a scalable, automated pipeline for labeling SWE-bench-style datasets with annotations for issue clarity, test coverage, and effort estimation. SPICE combines context-aware code navigation, rationale-driven prompting, and multi-pass consensus to produce labels that closely approximate expert annotations. SPICE’s design was informed by our own experience and frustration in labeling more than 800 instances from SWE-Gym. SPICE achieves strong agreement with human-labeled SWE-bench Verified data while reducing the cost of labeling 1,000 instances from around \$100,000 (manual annotation) to just \$5.10. These results demonstrate SPICE’s potential to enable cost-effective, large-scale dataset creation for SE-focused FMs. To support the community, we release both SPICE[Uncaptioned image]tool and SPICE[Uncaptioned image]Bench, a new dataset of 6,802 SPICE-labeled instances curated from 291 open-source projects in SWE-Gym (over 13x larger than SWE-bench Verified).
@@ I Introduction
In summary, manual labeling of SWE-bench-like datasets is time-consuming, expensive, subject to labeler bias, and does not scale. Towards tackling this challenge, we present SPICE[Uncaptioned image]: an automated platform for labeling SWE-bench-like datasets that strives for data quality at scale. Leveraging Aider [12], a context-aware code navigation and editing AI assistant, SPICE extracts relevant information from repositories, constructs structured prompts, and synthesizes high-quality labels for issue clarity and test coverage in a scalable manner. To enhance robustness, SPICE labels each instance multiple times and subsequently performs an aggregation through majority voting (a.k.a., self-consistency via stochastic sampling [13]). This process yields a more statistically grounded approximation of expert consensus.
@@
    Spice reduces the labeling cost of 1,000 instances from an estimated \$100,000 (using expert software engineer labelers) to just \$5.10.
@@ IV SPICE [Uncaptioned image]
IV-A Issue Clarity Assistance (ICA) Pipeline
@@ V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-V?
For $n=1,000,m=4,c=\$75,t=20min$, the central cost estimate is \$$\$100,000$ ($\eta_{c}=\eta_{t}=0$), the minimum cost $C_{\min}$ is \$$\$81,000$ ($\eta_{c}=-10\%,\eta_{t}=-10\%$) and the maximum cost $C_{\max}$ is \$$\$121,000$ ($\eta_{c}=10,\eta_{t}=10$). In other words, our projected cost for labeling 1,000 instances ranges from \$81k to \$121k, with a central value of \$100k. The estimate considers $n=4$, given that OpenAI’s-SWV [1] involved three labelers, and one moderator. While the moderator can be one of the labelers, they would still have invested cognition and time to reason through the other two labeler’s rationales to decide a final label, and note down her justification for her final choice. Hence, we assume that the average time of 20 minutes also applies for the moderator task as well.
@@
Most importantly, the total time taken by SPICE is approximately 53 times lower than that taken by humans: we estimate a total of 1,333 engineer-hours to label 1K instances (i.e., 20 min per instance * 4-expert panel * 1,000 instances), as opposed to 25 hours (i.e., approximately 90 sec per instance * 1000 instances) by SPICE.
@@
Table II provides the estimated labeling costs for 1,000 instances based on median token usage, should these models be used for both issue and test labeling. These costs combine total input and output tokens with their respective pricing structures. Remarkably, as per our default setting (SPICEG-4o-m for ICA and SPICEDS-R for TCA), labeling 1,000 instances using SPICE costs approximately \$5.1 (\$0.11 for issues and \$4.99 for tests). In contrast, our baseline cost using four senior software engineers is \$100,000. This translates to a remarkable cost reduction factor of approximately 19,600 times, highlighting SPICE’s extraordinary cost efficiency.
@@ VI Related Work
SWE-bench Limitations and Variants. While SWE-V (a subset of SWE-bench) largely superseded SWE-bench, it still has clear limitations. The dataset is relatively small (500 instances), all 12 projects are written in Python, and empirical evidence suggests that newer models were trained with SWE-bench data and have memorized issue-fix pairs [58] (SPICE can be used to create fresh datasets and mitigate the risks of benchmark data leakage during pre-training). Other research studies have recently flagged additional data quality issues in SWE-V [17].

### v3

@@ Abstract
High-quality labeled datasets are crucial for training and evaluating foundation models in software engineering, but creating them is often prohibitively expensive and labor-intensive. We introduce SPICE, a scalable, automated pipeline for labeling SWE-bench-style datasets with annotations for issue clarity, test coverage, and effort estimation. SPICE combines context-aware code navigation, rationale-driven prompting, and multi-pass consensus to produce labels that closely approximate expert annotations. SPICE’s design was informed by our own experience and frustration in labeling more than 800 instances from SWE-Gym. SPICE achieves strong agreement with human-labeled SWE-bench Verified data while reducing the cost of labeling 1,000 instances from around \$100,000 (manual annotation) to only \$5.10. These results demonstrate SPICE’s potential to enable cost-effective, large-scale dataset creation for SE-focused FMs. To support the community, we release both SPICE[Uncaptioned image]tool and SPICE[Uncaptioned image]Bench, a new dataset of 6,802 SPICE-labeled instances curated from 291 open-source projects in SWE-Gym (over 13x larger than SWE-bench Verified).
@@ I Introduction
In summary, manual labeling of SWE-bench-like datasets is time-consuming, expensive, subject to labeler bias, and does not scale. Compounding these challenges, the scarcity of high-quality, manually labeled benchmarks has driven the community to rely repeatedly on the same SWE-bench data—raising the risk of overfitting and potentially training on the ben hampering meaningful progress. Given that LLM-makers quite potentially train new LLMs on the benchmarking data [12], fresh benchmarks are required. To address these challenges, we present SPICE[Uncaptioned image]: an automated platform for labeling SWE-bench-like datasets that strives for data quality at scale. Leveraging Aider [13], a context-aware code navigation and editing AI assistant, SPICE extracts relevant information from repositories, constructs structured prompts, and synthesizes high-quality labels for issue clarity and test coverage in a scalable manner. To enhance robustness, SPICE labels each instance multiple times and subsequently performs an aggregation through majority voting (a.k.a., self-consistency via stochastic sampling [14]). This process yields a more statistically grounded approximation of expert consensus.
@@
    Spice reduces the labeling cost of 1,000 instances from an estimated \$100,000 (using expert software engineer labelers) to only \$5.10.
@@ IV SPICE [Uncaptioned image]
IV-A Issue Clarity Assessment (ICA) Pipeline
@@ V-B RQ2: How similar are the SPICE-generated rationales to those in SWE-V?
For $n=1,000,m=4,c=\$75,t=20min$, the central cost estimate is \$100,000\\$100,000\$ 100 , 000 ($\eta_{c}=\eta_{t}=0$), the minimum cost $C_{\min}$ is \$81,000\\$81,000\$ 81 , 000 ($\eta_{c}=-10\%,\eta_{t}=-10\%$) and the maximum cost $C_{\max}$ is \$121,000\\$121,000\$ 121 , 000 ($\eta_{c}=10,\eta_{t}=10$). In other words, our projected cost for labeling 1,000 instances ranges from \$81k to \$121k, with a central value of \$100k. The estimate considers $n=4$, given that OpenAI’s-SWV [1] involved three labelers, and one moderator. While the moderator can be one of the labelers, they would still have invested cognition and time to reason through the other two labeler’s rationales to decide a final label, and note down her justification for her final choice. Hence, we assume that the average time of 20 minutes also applies for the moderator task as well.
@@
Overall,the total time taken by SPICE is approximately 53 times lower than that taken by humans: we estimate a total of 1,333 engineer-hours to label 1K instances (i.e., 20 min per instance * 4-expert panel * 1,000 instances), as opposed to 25 hours (i.e., approximately 90 sec per instance * 1000 instances) by SPICE.
@@
Table II provides the estimated labeling costs for 1,000 instances based on median token usage, should these models be used for both issue and test labeling. These costs combine total input and output tokens with their respective pricing structures. As per our default setting (SPICEG-4o-m for ICA and SPICEDS-R for TCA), labeling 1,000 instances using SPICE costs approximately \$5.1 (\$0.11 for issues and \$4.99 for tests). In contrast, our baseline cost using four senior software engineers is \$100,000. This translates to a remarkable cost reduction factor of approximately 19,600 times, highlighting SPICE’s extraordinary cost efficiency.
@@ VI Related Work
SWE-bench Limitations and Variants. While SWE-V (a subset of SWE-bench) largely superseded SWE-bench, it still has clear limitations. The dataset is relatively small (500 instances), all 12 projects are written in Python, and empirical evidence suggests that newer models were trained with SWE-bench data and have memorized issue-fix pairs [59] (SPICE can be used to create fresh datasets and mitigate the risks of benchmark data leakage during pre-training). Other research studies have recently flagged additional data quality issues in SWE-V [18]. Importantly, SWE-V patches do not cover all corner cases, and incorrect patches can still sneak through [60].
@@ Acknowledgments
We would like to acknowledge the following people’s contributions to this work (ordered by last name): Yuan Guowen, Hu Jiajun, and Chong Chun Yong.